OpenStudy (anonymous):

solve. x-1 over 4 minus 3x+2 over 6=1/2

OpenStudy (anonymous):

x=.5 repeating

OpenStudy (anonymous):

x=-13/3

OpenStudy (anonymous):

how did u get x is -13/3

OpenStudy (anonymous):

Multiply both sides by 24: 6(x-1)-4(3x+2)=12 6x-6-12x-8=12 -6x=26 x=26/-6

OpenStudy (anonymous):

=-13/3

OpenStudy (anonymous):

Thanks blender

OpenStudy (anonymous):

Np :)

OpenStudy (anonymous):

The first one (.555) was adding the two instead of subtracting :X